If the box accelerated at 0.5 m/s^2 to the right
Answer #1

fk = 6 N
M = 40 Kg
angle = 30 degree

Let the force be F
Hence,
Fx = Fcos30
Fy = Fsin30

a) a = 0.5 m/s^2

Fx - fk = ma
Fcos30 - 6 = 40*0.5
Fcos30 = 6 + 20
Fcos30 = 26
F = 30.022

Fx = Fcos30 = 26 N
Fy = Fsin30 = 15.011 N

b) F = 30.022 N
c) Ground Exerts a Force = mg - Fy
= 40*9.8 - 15.011
= 376.989 N

d) 6 = muk*m*g
6 = muk*40*9.8
muk = 6/(40*9.8)
muk = 0.01531